# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ Abstract
The primordial neutrino asymmetry leave profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS). We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry $\xi_{\nu}$, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism. The effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$. Then we perform the state-of-the-art calculation for the abundance of light elements including (anti)neutrino spectral distortions, which indicate a positive asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ from EMPRESS data. The implications of the neutrino asymmetry for the CMB and LSS are studied in detail, and we find that the Baryon Acoustic Oscillations (BAO) are also significantly affected by $\xi_{\nu}$ in addition to the sum of neutrino masses. A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.013$, which provides constraints on UV models capable of producing large asymmetries.
@@ 1 Introduction
A precise understanding of the neutrino decoupling process is essential for accurate constraints on neutrino properties. Notably, the partial reheating of neutrinos and antineutrinos due to the small overlap between neutrino decoupling and $e^{\pm}$ annihilations results in slightly non-thermal spectra and an increased total neutrino energy density, typically described by the effective number of relativistic species, $N_{\rm eff}$. The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes. For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv . Including flavor oscillations, matter effects, finite-temperature QED corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}(0)=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
@@
On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention. Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons, and photons, respectively. However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
@@
The lepton asymmetries prior to the neutrino decoupling epoch, hereafter referred to as primordial neutrino asymmetries, have significant implications for the evolution of the universe. During the neutrino decoupling epoch, primordial neutrino asymmetries lead to an increased $N_{\rm eff}$ and specific spectral distortions of neutrinos and antineutrinos Grohs:2016cuu . Furthermore, if the asymmetries among different flavors are different, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , resulting in the redistribution of asymmetries among flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz . After neutrino decoupling, during the BBN epoch, primordial neutrino asymmetries primarily affect the final positron-to-neutron ratio and thus the final abundances of light elements. This effect mainly depends on the density distributions of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are subleading Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx . Additionally, primordial neutrino asymmetries influence CMB anisotropies and LSS through increased $N_{\rm eff}$ and modified light element abundances, alongside the effects of massive neutrinos. Consequently, observations of BBN, CMB, and LSS can be utilized to constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . In particular, recent studies from the EMPRESS survey indicate a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ represents the degeneracy parameter for electron neutrinos.
@@
However, the current situation is less satisfactory compared to the Standard Model (SM) case due to persistent ambiguities regarding neutrino decoupling. These ambiguities particularly involve neutrino spectral distortions and the effects of flavor oscillations and QED corrections. As a result, the effects of primordial neutrino asymmetries on BBN and CMB, as well as the constraints derived from observational data, are generally evaluated using approximations. These approximations often include the instantaneous neutrino decoupling approximation and the assumption of massless neutrinos. To derive precise and accurate constraints on primordial neutrino asymmetries from both current and future observations, it is essential to rigorously analyze the implications of these asymmetries through a comprehensive treatment of the neutrino decoupling process. A recent study Froustey:2024mgf makes effort toward a rigorous analysis of primordial neutrino asymmetries. This study focuses on the implications of neutrino synchronous oscillations for BBN, while ignoring the correction from (anti)neutrino spectral distortions, etc, and furthermore, the implications for both the CMB and LSS are not addressed at all.
@@
In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB and LSS in the presence of primordial neutrino asymmetries. To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, it is crucial to employ the complete set of neutrino quantum kinetic equations (QKEs). These equations can be derived using various approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and more generally from the Closed-Time Path (CTP) formalism of non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt . The main improvements in our study are as follows:
@@
  • We numerically solve the QKEs for the neutrino decoupling process in the presence of primordial neutrino asymmetry, including effects of flavor oscillations, matter interactions, finite-temperature QED (FTQED) corrections up to $\mathcal{O}(3)$ order, and complete neutrino-electron and neutrino-neutrino collision terms. This approach matches the accuracy used in current state-of-the-art predictions for the Standard Model (SM) case Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , are still debated for the SM case and are not included in this study..
  • We use the resulting density matrices to study the impacts of primordial neutrino asymmetry on the BBN, CMB, and LSS. The effects of neutrino spectral distortions are also highlighted by comparing these results with those obtained using the Fermi-Dirac distribution for neutrinos.
  • With the observation that the Baryon Acoustic Oscillations (BAO) are also affected by the primordial neutrino asymmetry, we derive the constraints on the primordial neutrino asymmetry by combining the traditional BBN and CMB datasets with the BAO observations for the first time.
Figure 1: A schematic illustration outlining the structure of this paper. Section 2 presents the derivation of the QKEs for neutrinos within the CTP formalism. These QKEs are subsequently applied to study neutrino decoupling in the presence of primordial neutrino asymmetry $\xi_{\nu}$ in Section 3. Based on such precise treatment of neutrino decoupling, Sections 4 and 5 explore the effects of primordial neutrino asymmetry on BBN, CMB and LSS. For each section, the primary input and output physical quantities that depend on $\xi_{\nu}$ are labelled in each box. Finally, in Section 6, cosmological constraints on $\xi_{\nu}$ are obtained by the MCMC analyses combining the theoretical predictions and observational data from BBN, CMB, and LSS.
@@
  • We numerically evaluate $N_{\rm eff}$ and the (anti)neutrino spectral distortions, incorporating primordial neutrino asymmetries with state-of-the-art accuracy. We find that $N_{\rm eff}$ can be expressed as $N_{\rm eff}=N_{\rm eff}^{\rm SM}+\delta N_{\rm eff}^{\xi_{\nu},\rm ther}+0.0102\,\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.002$ represents the SM prediction with comparable precision Froustey:2020mcq ; Bennett:2020zkv . The term $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}$ accounts for the contribution from primordial neutrino asymmetry under the instantaneous decoupling approximation, and is given by $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}\simeq 3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$, while the final term represents the additional correction from non-instantaneous decoupling effects in the presence with non-zero primordial neutrino asymmetry.
@@
  • We provide state-of-the-art predictions for light element abundances produced by BBN, through which the EMPRESS measurement on helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$ immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$. We also find that the (anti)neutrino spectral distortions are important for precise predictions for BBN, especially for the abundance of Helium-4 $Y_{P}$.
@@
  • For the constraints on primordial neutrino asymmetry, combing EMPRESS BBN and Planck CMB data yields $\xi_{\nu}=0.028\pm 0.013$, while incorporating BOSS BAO data in the analysis yields $\xi_{\nu}=0.024\pm 0.013$. Compared to previous result obtained in ref $\xi_{\nu}=0.034\pm 0.014$, our results indicate a $\sim 18\%$ and $\sim 29\%$ smaller central value and a slight reduction in the uncertainty, respectively.
@@
  • The impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them are also discussed. Finally, we show how the constraints on $\xi_{\nu}$ can be converted into the constraints on UV model parameters with a specific Q-ball decay model capable of producing large primordial neutrino asymmetries.
@@
This paper is organized as follows. In Section 2, we re-derive the QKEs for Dirac and Majorana neutrinos using the CTP formalism. In Section 3, we present the QKEs in the context of the early universe and provides corresponding numerical results considering primordial neutrino asymmetries. The impacts of primordial neutrino asymmetries on the BBN epoch is discussed in Section 4, while their effects on both CMB and LSS are analyzed in Section 5. In Section 6, we present the results from a joint MCMC analysis of neutrino primordial asymmetries and neutrino masses, utilizing datasets from BBN, CMB and BAO observations. Finally, we conclude in Section 7. To guide the readers, a schematic illustration of this work are presented in Fig. 1. Throughout this manuscript, we use the natural units where $\hbar=c=k_{B}=1$.

### v2

@@ Abstract
The primordial neutrino asymmetry leave profound imprints on the evolution history of the universe, which can be constrained by cosmological observations, including Big Bang Nucleosynthesis (BBN), Cosmic Microwave Background (CMB), and Large-Scale Structure (LSS). We present comprehensive analysis on implications and constraints of the primordial neutrino asymmetry $\xi_{\nu}$, based on a precise treatment of neutrino decoupling by solving the complete (anti)neutrino quantum kinetic equations in the Closed-Time-Path formalism. The effective number of neutrinos $N_{\rm eff}$ and (anti)neutrino spectral distortions are calculated, and we find that the non-instantaneous decoupling correction is given by $\delta N_{\rm eff}=0.0440+0.0102\,\xi_{\nu}^{2}$. Then we perform the state-of-the-art calculation for the abundance of light elements including (anti)neutrino spectral distortions, which indicate a positive asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ from EMPRESS data. The implications of the neutrino asymmetry for the CMB and LSS are studied in detail, and we find that the Baryon Acoustic Oscillations (BAO) are also significantly affected by $\xi_{\nu}$ in addition to the sum of neutrino masses. A combined analysis with EMPRESS BBN, Planck CMB and BOSS BAO data yields a tighter constraint $\xi_{\nu}=0.024\pm 0.012$, which provides constraints on UV models capable of producing large asymmetries.
@@ 1 Introduction
A precise understanding of the neutrino decoupling process is essential for accurate constraints on neutrino properties. Notably, the partial reheating of neutrinos and antineutrinos due to the small overlap between neutrino decoupling and $e^{\pm}$ annihilations results in slightly non-thermal spectra and an increased total neutrino energy density, typically described by the effective number of relativistic species, $N_{\rm eff}$. The values of $N_{\rm eff}$ and the spectral distortions of neutrinos and antineutrinos are used as initial conditions to evaluate the impact of relic neutrinos in subsequent processes. For the Standard Model case, where asymmetries between neutrinos and antineutrinos are assumed to be negligible, significant progress has been made in detailing the neutrino decoupling process over recent decades Hannestad:1995rs ; Dolgov:1997mb ; Mangano:2001iu ; Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq ; Akita:2020szl ; Froustey:2019owm ; EscuderoAbenza:2020cmq ; Froustey:2020mcq ; Bennett:2020zkv . Including flavor oscillations, matter effects, Finite-Temperature QED (FTQED) corrections up to $\mathcal{O}(3)$, and full neutrino-electron and neutrino-neutrino collision terms, the current standard prediction yields $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv ; Akita:2020szl .
@@
On the other hand, inspired by the recent EMPRESS survey on primordial helium abundance empress and its potential to alleviate cosmological tensions Barenboim:2016lxv ; Yeung:2020zde ; Seto:2021tad ; Kumar:2022vee ; Yeung:2024krv , the scenario where asymmetries between neutrinos and antineutrinos are not negligible has recently attracted significant attention. Due to the sphaleron processes in the early universe Kuzmin:1985mm ; Khlebnikov:1988sr ; Harvey:1990qw ; Dreiner:1992vm , one might naively expect neutrino asymmetries, or lepton asymmetries, to be of the same order as the baryon asymmetry of the universe, which is strongly constrained by BBN and CMB observations: $\eta_{B}\equiv(n_{B}-n_{\bar{B}})/n_{\gamma}=(6.14\pm 0.04)\times 10^{-10}$ Planck:2018vyg , where $n_{B}$, $n_{\bar{B}}$, and $n_{\gamma}$ denote the number densities of baryons, antibaryons and photons, respectively. However, substantial lepton asymmetries compared to the baryon asymmetry before the neutrino decoupling epoch can still be generated by various models, see $e.g.,$ Casas:1997gx ; Dolgov:1989us ; Bajc:1997ky ; Asaka:2005pn ; Asaka:2005an ; Pilaftsis:2003gt ; Borah:2022uos ; Kawasaki:2002hq ; Kawasaki:2022hvx .
@@
The lepton asymmetries present before the neutrino decoupling epoch, henceforth referred to as primordial neutrino asymmetries, have significant implications for the universe’s evolution. During the neutrino decoupling epoch, these asymmetries contribute to an increased $N_{\rm eff}$ and induce specific spectral distortions in neutrinos and antineutrinos Grohs:2016cuu . If the asymmetries vary between different neutrino flavors, neutrinos and antineutrinos can undergo synchronous oscillations Bell:1998ds ; Pastor:2001iu ; Abazajian:2002qx ; Wong:2002fa , leading to a redistribution of asymmetries among the flavors Dolgov:2002ab ; Pastor:2008ti ; Mangano:2010ei ; Mangano:2011ip ; Castorina:2012md ; Barenboim:2016shh ; Johns:2016enc ; Froustey:2021azz . After neutrino decoupling, during the Big Bang Nucleosynthesis (BBN) epoch, primordial neutrino asymmetries primarily affect the final proton-to-neutron ratio and the resulting light element abundances. This effect depends largely on the distribution of electron neutrinos and antineutrinos, while corrections from the increased $N_{\rm eff}$ are secondary Sarkar:1995dd ; Iocco:2008va ; Pitrou:2018cgg ; Serpico:2005bc ; Mangano:2011ip ; Chu:2006ua ; Simha:2008mt ; Saviano:2013ktj ; Burns:2023sgx . Additionally, these asymmetries influence cosmic microwave background (CMB) anisotropies and large-scale structure (LSS) through the increased $N_{\rm eff}$ and modified light element abundances, along with the effects of massive neutrinos. Consequently, observations of BBN, CMB, and LSS can constrain primordial neutrino asymmetries Simha:2008mt ; Oldengott:2017tzj ; Pitrou:2018cgg ; Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okz . Notably, recent findings from the EMPRESS survey suggest a preference for a positive asymmetry in electron neutrinos Matsumoto:2022tlr ; Burns:2022hkq ; Escudero:2022okzx. Specifically, combining BBN and CMB observations yields $\xi_{\nu_{e}}=0.034\pm 0.014$ Escudero:2022okz , where $\xi_{\nu_{e}}=\mu_{\nu_{e}}/T_{\nu}$ is the degeneracy parameter for electron neutrinos.
@@
However, the current situation is less satisfactory compared to the Standard Model (SM) scenario due to persistent uncertainties concerning neutrino decoupling. These uncertainties involve neutrino spectral distortions, flavor oscillations, and finite-temperature quantum electrodynamics (FTQED) corrections. Consequently, the effects of primordial neutrino asymmetries on BBN and CMB, as well as the constraints derived from observational data, are generally evaluated using approximations, such as the instantaneous neutrino decoupling approximation and the assumption of massless neutrinos. To derive precise constraints on primordial neutrino asymmetries from both current and future observations, it is essential to rigorously analyze these asymmetries within a comprehensive framework of neutrino decoupling. A recent study Froustey:2024mgf makes an effort towards a rigorous analysis of primordial neutrino asymmetries. This study focuses on the implications of neutrino synchronous oscillations on the neutrino flavor equilibration and BBN process, while the corresponding constraints on the primordial neutrino asymmetries are obtained by combining the BBN measurements and the constraints on the parameter set ($N_{\rm eff},Y_{P},\omega_{b}$) from CMB + BAO experiments. However, the resulting (anti)neutrino spectral distortions and their implications for the BBN and CMB are not discussed in Froustey:2024mgf . Furthermore, the implications of primordial neutrino asymmetries for both the CMB angular power spectrum and the BAO are not addressed in Froustey:2024mgf , which according to our results are actually important for the accurate determination of the cosmological constraints.
@@
In this work, we focus on evaluating the neutrino decoupling process and its subsequent effects on BBN, CMB, and LSS in the presence of primordial neutrino asymmetries. To accurately describe the evolution of neutrino density matrices during decoupling, including flavor oscillations and full collision effects, we employ the complete set of neutrino quantum kinetic equations (QKEs). These equations are derived using several approaches, such as perturbative expansions of the density matrix Sigl:1993ctk , relativistic generalizations of the Bogoliubov-Born-Green-Kirkwood-Yvon (BBGKY) equations Volpe:2013uxl ; Volpe:2015rla ; Froustey:2020mcq , and the Closed-Time Path (CTP) formalism from non-equilibrium quantum field theory Vlasenko:2013fja ; Blaschke:2016xxt ; Kainulainen:2023ocv ; Kainulainen:2024fdg . The main improvements of our study are as follows:
@@
  • We numerically solve the QKEs for the neutrino decoupling process, accounting for flavor oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and complete neutrino-electron and neutrino-neutrino collision terms, matching the accuracy of state-of-the-art Standard Model (SM) predictions Froustey:2020mcq ; Bennett:2020zkv  [^1] Additional QED corrections, such as those to interaction vertices Cielo:2023bqp ; Jackson:2023zkl ; Drewes:2024wbw , remain under debate for the SM case and are not included in this study..
  • We use the resulting density matrices to study the impacts of primordial neutrino asymmetry on BBN, CMB, and LSS, with a particular emphasis on (anti)neutrino spectral distortions by comparing these results with those obtained using the Fermi-Dirac distribution for neutrinos.
  • Recognizing that Baryon Acoustic Oscillations (BAO) are also influenced by primordial neutrino asymmetry, we derive constraints on primordial neutrino asymmetry by combining traditional BBN and CMB datasets with BAO observations for the first time.
Figure 1: A schematic illustration of the paper’s structure. Section 2 derives the QKEs for neutrinos using the CTP formalism. These QKEs are then applied to study neutrino decoupling in the presence of primordial neutrino asymmetry $\xi_{\nu}$ in Section 3. Based on this precise treatment of neutrino decoupling, Sections 4 and 5 explore the effects of primordial neutrino asymmetry on BBN, the CMB, and LSS. For each section, the primary input and output physical quantities dependent on $\xi_{\nu}$ are labeled in each box. Finally, in Section 6, we obtain cosmological constraints on $\xi_{\nu}$ through Markov Chain Monte Carlo (MCMC) analyses, combining theoretical predictions with observational data from BBN, CMB, and LSS.
@@
  • We numerically evaluate $N_{\rm eff}$ and the (anti)neutrino spectral distortions, incorporating primordial neutrino asymmetries with state-of-the-art accuracy. For primordial neutrino asymmetries with complete flavor equilibration before neutrino decoupling, $i.e.,$ $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we find that $N_{\rm eff}$ can be expressed as $N_{\rm eff}=N_{\rm eff}^{\rm SM}+\delta N_{\rm eff}^{\xi_{\nu},\rm ther}+0.0102\,\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.002$ represents the SM prediction Froustey:2020mcq ; Bennett:2020zkv . The term $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}$ accounts for the contribution from primordial neutrino asymmetry under the instantaneous decoupling approximation and is given by $\delta N_{\rm eff}^{\xi_{\nu},\rm ther}\simeq 3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)$. The final term represents the additional correction from non-instantaneous decoupling effects in the presence of non-zero primordial neutrino asymmetry.
@@
  • We provide state-of-the-art predictions for the abundances of light elements produced by BBN. Using the EMPRESS measurement of helium abundance, $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, we immediately indicate a positive primordial neutrino asymmetry, $0.032\leq\xi_{\nu}\leq 0.052$, with fixed baryon abundance $\omega_{b}$. We also find that (anti)neutrino spectral distortions are crucial for precise BBN predictions, especially for the abundance of Helium-4, $Y_{P}$.
@@
  • For constraints on primordial neutrino asymmetry, combining EMPRESS BBN and Planck CMB data yields $\xi_{\nu}=0.028\pm 0.012$, while incorporating BOSS BAO data results in $\xi_{\nu}=0.024\pm 0.012$. Compared to the previous result obtained in Ref. Escudero:2022okz , $\xi_{\nu}=0.034\pm 0.014$, our results show an approximately 18% and 29% reduction in the central value and a slight decrease in uncertainty, respectively.
@@
  • We also discuss the impacts of treating both the sum of neutrino masses $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each. Finally, we demonstrate how the constraints on $\xi_{\nu}$ can be converted into constraints on ultraviolet (UV) model parameters, using a specific Q-ball decay model capable of generating large primordial neutrino asymmetries.
@@
This paper is organized as follows. In Section 2, we re-derive the QKEs for Dirac and Majorana neutrinos using the CTP formalism. Section 3 presents the QKEs in the context of the early universe and provides corresponding numerical results, considering primordial neutrino asymmetries. The impact of primordial neutrino asymmetries on the BBN epoch is discussed in Section 4, while their effects on both the CMB and LSS are analyzed in Section 5. In Section 6, we present results from a joint MCMC analysis of neutrino primordial asymmetries and neutrino masses, utilizing datasets from BBN, CMB, and BAO observations. Finally, we conclude in Section 7. To guide the reader, a schematic illustration of this work is presented in Fig. 1. Throughout this manuscript, we use natural units where $\hbar=c=k_{B}=1$.